# In-Context Learning Enhanced Credibility Transformer

### v1

@@ Abstract
The starting point of our network architecture is the Credibility Transformer which extends the classical Transformer architecture by a credibility mechanism to improve model learning and predictive performance. This Credibility Transformer learns credibilitized CLS tokens that serve as learned representations of the original input features. In this paper we present a new paradigm that augments this architecture by an in-context learning mechanism, i.e., we increase the information set by a context batch consisting of similar instances. This allows the model to enhance the CLS token representations of the instances by additional in-context information and fine-tuning. We empirically verify that this in-context learning enhances predictive accuracy by adapting to similar risk patterns. Moreover, this in-context learning also allows the model to generalize to new instances which, e.g., have feature levels in the categorical covariates that have not been present when the model was trained – for a relevant example, think of a new vehicle model which has just been developed by a car manufacturer.
@@ 1 Introduction
A fundamental challenge in actuarial modeling lies in balancing predictive performance, explainability and reliable model estimation. This is especially true for risks for which one does not have a long observation history. Classical credibility theory, formalized by Bühlmann, (1967) and Bühlmann and Straub, (1970), provides an elegant framework for addressing this challenge through an optimal linear combination of instance individual and collective claims experience. In parallel, the machine learning community has developed sophisticated approaches for learning from limited data, culminating in the emergence of in-context learning (ICL) capabilities in large-scale Transformer models. The discovery that large-scale Transformers can adapt to new tasks through demonstration examples, revolutionized our understanding of few-shot learning Brown et al., (2020). That is, based on off-line training, a large-scale Transformer architecture can perform new tasks if sufficient context to these new tasks is provided, in particular, these tasks are done without retraining the Transformer architecture to the new situation. In some sense, this resembles the capability of being able to perform transfer learning. Recent theoretical work has established that this phenomenon can be understood as implicit gradient descent Akyürek et al., (2023), Bayesian inference Xie et al., (2022) and ICL on tabular data Hollmann et al., (2023) and Müller et al., (2024).
@@
Recent work by Richman et al., 2025a introduced the Credibility Transformer, demonstrating how classical credibility concepts can be embedded within modern Transformer architectures helping to improve the predictive performance of Transformers on tabular input data. Their approach leverages the CLS token mechanism of Devlin et al., (2019) to implement a learnable prior that is combined with feature-specific information through an attention-weighted averaging – similarly to linear Bühlmann, (1967) credibility – achieving state-of-the-art performance on insurance datasets while maintaining explainability. A main benefit of this credibility mechanism that acts similar to drop-out during model training is the stabilization of the training algorithm leading to delayed early stopping points and better feature extraction.
@@
However, existing approaches, including the Credibility Transformer, evaluate each risk in isolation during inference, neglecting potentially valuable information from similar observations within the portfolio. This limitation becomes particularly relevant in practical scenarios where insurers must price policies for risk profiles with limited historical experience while having access to recent claims data from similar risks, just think of expanding an existing insurance product to new geographical regions. The ability to leverage this contextual information dynamically, without retraining the model, may significantly enhance predictive accuracy and adapt to risk patterns of similar experience.
@@ 1.1.1 Classical credibility theory
Credibility theory provides the theoretical foundation for combining individual experience with collective information in actuarial applications. The seminal work of Bühlmann, (1967) and Bühlmann and Straub, (1970) established the mathematical framework for an optimal linear prediction taking into account both sources of information. Generally, combining different sources of information involves intractable Bayesian posterior computations, and coping with this intractability, Bühlmann, (1967) proposed a best-linear approximation by minimizing the mean squared error (MSE). This linear approximation takes the form
@@
where $Y$ reflects the individual experience, $\mu_{0}$ is the collective information, and $\alpha\in[0,1]$ is the credibility weight that weighs these two components. The credibility weight takes the form
$\alpha=\frac{w}{w+\kappa},$
where $w>0$ is a case weight indicating the amount of information contained in the individual observation $Y$ and $\kappa>0$ is the credibility coefficient (hyper-parameter) that balances between individual and collective information. Basically, this credibility coefficient gives the trade-off between the different uncertainties involved at the individual and the collective level; see Bühlmann and Gisler, (2005) for an extended discussion.
@@ 1.1.2 Transformer architectures for tabular data
The adaptation of the Transformer architecture of Vaswani et al., (2017) to tabular input data presents some challenges compared to their applications in large language models. Unlike sequential or grid-structured data, tabular features lack an inherent time-causal relationship for a canonical tensor encoding. This requires novel approaches to tokenization and positional encoding. The TabTransformer of Huang et al., (2020) pioneered the application of Transformers to tabular data by applying self-attention exclusively to categorical embeddings while processing continuous features through a separate pathway. This design choice reflects the observation that categorical variables often exhibit complex interactions that benefit from attention mechanisms, while continuous features may adequately be handled by simpler feed-forward neural network (FNN) layers.
The feature tokenizer (FT) Transformer of Gorishniy et al., (2021) extends this approach by tokenizing all features, both categorical and continuous ones, and processing them through a unified Transformer architecture. Their continuous feature tokenization uses a simple FNN embedding with an embedding space dimension being identical to the one chosen for the categorical entity embedding. This then naturally allows one to write the input feature in tensor form, which is the suitable structure for Transformer architectures. In an actuarial context, Brauer, (2024) has empirically verified the excellent performance of these TF Transformer architectures on tabular input data.
Recent work on numerical feature encoding has explored various alternative approaches to handle continuous variables in Transformers. Noteworthy, piecewise linear encoding (PLE) of Gorishniy et al., (2022) provides a differentiable binning of continuous features that adapts to the data distribution. Periodic encoding using sinusoidal transformations Tancik et al., (2020) can capture cyclical patterns (maybe less relevant in general insurance pricing), while quantile-based encoding leverages distributional information for a robust feature representation.
@@ 1.1.3 In-context learning theory
ICL emerged as a surprising capability of large language models, where models can adapt to new tasks through demonstration examples provided in the input, without any parameter updates and model retraining; see Brown et al., (2020). This phenomenon has sparked intensive research into understanding the underlying mechanisms and theoretical foundations of ICL. Akyürek et al., (2023) demonstrated that linear self-attention layers can implement gradient descent on a MSE objective function when provided with appropriate in-context examples. Their analysis shows that Transformers can internally implement optimization algorithms, with attention weights encoding gradient information and value projections performing parameter updates. This finding suggests that ICL implements a form of implicit meta-learning within the forward pass. From a Bayesian perspective, Xie et al., (2022) and Müller et al., (2024) interpret ICL as implicit Bayesian inference, where the model maintains a posterior distribution over possible tasks and updates this posterior based on the context examples. Under this framework, the pre-trained distribution defines the prior, and attention mechanisms implement approximate posterior inference, providing a probabilistic foundation for understanding ICL behavior. This interpretation closes the loop to Bühlmann, (1967), and it connects classical actuarial thinking with modern machine learning concepts.
Recent work has identified important factors that influence ICL performance. Min et al., (2022) demonstrate that the format and distribution of demonstration examples matter more than their correctness, suggesting that ICL relies heavily on pattern matching and format recognition, while Wei et al., (2023) show that ICL capabilities emerge at scale, with larger models exhibiting qualitatively different learning behaviors than smaller ones.
Within the tabular learning context, where models are usually trained until performance on a validation loss is optimal, it would be surprising to find that ICL could lead to additional significant improvements in performance and – if this were to be the case – a new explanation for this improvement would be needed. Whereas Müller et al., (2024) have indeed shown that their TabICL model outperforms many other models, including models that only use supervised learning, as well as those using ICL, nonetheless, this could be explained by their extensive foundation model pre-training scheme using real and synthetic data. Classical supervised learning models are typically trained on a finite sample using early stopping of the gradient descent algorithm to prevent over-fitting. This early stopping may also prevent learning a more detailed, fine-grained structure, and foundation models being trained on a large body of (similar) data may be aware of such a detailed structure.
In this work, we would like to study the following questions: First, can ICL improve performance on a model that has been pre-trained using supervised learning? Second, if this is the case, what is the explanation for this improvement? Third, can ICL for tabular data be used to improve the performance of a much smaller pre-trained model focussing only on a single dataset?
@@ 1.1.4 Connections between credibility and meta-learning
The connection between actuarial credibility theory and modern meta-learning approaches has received limited attention despite their conceptual similarities. Both frameworks address the fundamental challenge of making predictions with limited task-specific data by leveraging broader experience. In credibility theory, this manifests as combining individual risk experience with portfolio-level information. In meta-learning, it involves leveraging experience across multiple tasks to enable rapid adaptation to new tasks. The mathematical structures underlying these approaches exhibit close parallels. The credibility formula in (1.1) can be viewed as a special case of prototype-based meta-learning, where the collective mean $\mu_{0}$ serves as a prototype and the credibility factor $\alpha$ determines the adaptation strength. Similarly, the attention mechanisms in Transformers implement a form of non-parametric regression that generalizes credibility weighting to high-dimensional feature spaces. Recent work in neural processes Garnelo et al., (2018) provides a framework that unifies these perspectives. Neural processes implement a form of meta-learning that closely resembles Bayesian inference, with the context set defining a posterior distribution over functions. This framework naturally accommodates varying amounts of context information, analogous to how credibility factors adjust (based on available experience) through the case weights $w>0$.
@@ 1.2 In-context learning enhanced Credibility Transformer
Our main contribution is the introduction of the In-Context Learning enhanced Credibility Transformer (ICL-Credibility Transformer), which combines the methods discussed in the previous subsections. First, the Credibility Transformer of Richman et al., 2025a leverages the FT Transformer with a credibility mechanism that significantly enhances model training. A crucial ingredient of the Credibility Transformer is the inclusion of the CLS token of Devlin et al., (2019), that is used to learn an encoding of the input information using the classical attention layer mechanism of Vaswani et al., (2017). But instead of simply encoding and extracting this CLS token, it is credibilitized with a linear credibility mechanism that combines the feature individual CLS token encoding with a collective information token. This linear credibility combination has a very positive effect on gradient descent training, resulting in excellent predictive models. In fact, during gradient descent training, a randomized version of the credibility mechanism is applied that either selects the instance-specific CLS token or the collective information token. This randomized mechanism during model training can be interpreted as a drop-out mechanism, similar to Srivastava et al., (2014) and Wager et al., (2013), but instead of just dropping out and masking, at the same time a collective information representation is learned on the dropped out instances. This construction results in the credibilitized CLS token which is a robustified encoding (representation) of the input data. The Credibility Transformer then processes these tokens through a decoder forming the predictions.
Inspired by the ICL architectures of Hollmann et al., (2023) and Müller et al., (2024) on tabular data, we modify the Credibility Transformer architecture to facilitate ICL. The crucial questions we were facing are:
  • (1)
@@
The primary contributions of this paper are:
@@
    We use the same projection layer as in the base Credibility Transformer to decode the outcome decorated CLS tokens. This is justified by the fact mentioned in item (b) saying that the ICL takes place in the space of the CLS tokens.
@@
We establish formal connections between ICL mechanisms and classical credibility theory, proving that our ICL basically represents a credibility mechanism that regulates the impact of the context on the target instances to be predicted. We demonstrate that the attention weights serve as adaptive, data-driven credibility factors that depend on both feature similarity and exposure levels, generalizing the fixed credibility weights of classical approaches.
@@
We deploy a two-phase training strategy that first establishes strong feature representations through a standard supervised learning method of the base Credibility Transformer. This learned base Credibility Transformer already presents a strong predictive model. The second training phase fine-tunes the model using ICL. During this phase, we freeze the decoder layers (i.e., the output decoder) and train the ICL mechanism, as well as the base Credibility Transformer encoder. This constrained training approach enhances the robustness of the learning process, while allowing the weights of the base Credibility Transformer encoder which were initially optimized for supervised learning to be modified. Optionally, a third training step can be applied by jointly fine-tuning all network components, i.e., including the decoder layer; however, this requires careful regularization and a small learning rate to prevent performance degradation.
@@
We verify on the popular French Motor Third Party Liability (MTPL) dataset of Dutang et al., (2024) the outstanding performance of the ICL-Credibility Transformer. Our study quantifies the contribution of each architectural component, while explainability analyses reveal that the learned attention patterns align with actuarial intuition about risk similarity and credibility weighting. An interesting property of ICL is that we can explicitly illustrate and try to understand how individual instances in the context batch enhance the predictions; we provide empirical results explaining this case by case. A second interesting case study shows how the ICL-Credibility Transformer can be used to predict on data with new covariate levels, e.g., a new car brand that has not been available during model training. This generalization property (similar to transfer learning) especially draws on the context, to find the right level of integration into the existing model.

### v2

@@ Abstract
The starting point of our network architecture is the Credibility Transformer which extends the classical Transformer architecture by a credibility mechanism to improve model learning and predictive performance. This Credibility Transformer learns credibilitized CLS tokens that serve as learned representations of the original input features. In this paper, we present a new paradigm that augments this architecture by an in-context learning mechanism, i.e., we increase the information set by a context batch of similar instances. This allows the model to enhance the CLS token representations of the instances by additional in-context information and fine-tuning. We empirically verify that this in-context learning improves predictive performance by adapting to similar risk patterns. Moreover, this in-context learning also allows the model to generalize to new instances which, e.g., have feature levels in the categorical covariates that have not been present when the model was trained – for a relevant example, think of a new vehicle model which has just been launched by a car manufacturer.
@@ 1 Introduction
A fundamental challenge in actuarial modeling lies in balancing predictive performance, explainability and reliable model estimation. This is especially true for risks for which one does not have a long observation history. Classical credibility theory, formalized by Bühlmann (1967) and Bühlmann and Straub (1970), provides an elegant framework for addressing this challenge through an optimal linear combination of instance individual and collective claims experience. In parallel, the machine learning community has developed sophisticated approaches for learning from limited data, culminating in the emergence of in-context learning (ICL) capabilities in large-scale Transformer models. The discovery that large-scale Transformers can adapt to new tasks through demonstration examples, revolutionized our understanding of few-shot learning Brown et al. (2020). That is, based on off-line training, a large-scale Transformer architecture can perform new tasks if sufficient context to these new tasks is provided. In particular, these tasks are done without retraining the Transformer architecture to the new situation. In some sense, this resembles the capability of being able to perform transfer learning. Recent theoretical work has established that this phenomenon can be understood as implicit gradient descent Akyürek et al. (2023) or Bayesian inference Xie et al. (2022). For ICL on tabular data, see Hollmann et al. (2023) and Müller et al. (2024).
@@
Recent work by Richman et al. (2025a) introduced the Credibility Transformer, demonstrating how classical credibility concepts can be embedded within modern Transformer architectures helping to improve the predictive performance of Transformers on tabular input data. Their approach leverages the classify (CLS) token mechanism of Devlin et al. (2019) to implement a learnable prior that is combined with feature-specific information through an attention-weighted averaging – similarly to linear Bühlmann (1967) credibility. This Credibility Transformer architecture achieves state-of-the-art performance on insurance datasets while maintaining some explainability. A main benefit of the credibility mechanism (that acts similar to drop-out during model training) is the stabilization of the training algorithm leading to a delayed early stopping point and better feature extraction.
@@
However, existing approaches, including the Credibility Transformer, evaluate each risk in isolation during inference, neglecting potentially valuable information from similar instances within the portfolio. This limitation becomes particularly relevant in practical scenarios where insurers must price policies for risk profiles with a limited historical experience while having access to recent claims data from similar policies, just think of expanding an existing insurance product to a new geographical region. The ability to leverage this contextual information dynamically, without retraining the model, may significantly enhance predictive accuracy and adapt to risk patterns of similar experience.
@@ 1.1.1 Classical credibility theory
Credibility theory provides the theoretical foundation for combining individual experience with collective information in actuarial applications. The seminal work of Bühlmann (1967) and Bühlmann and Straub (1970) established the mathematical framework for an optimal linear prediction taking into account both sources of information. In general, combining different sources of information involves intractable Bayesian posterior computations, and coping with this intractability, Bühlmann (1967) proposed a best-linear approximation by minimizing the mean squared error (MSE). This linear approximation takes the form
@@
where $Y$ reflects the individual experience, $\mu_{0}$ is the collective information, and $\alpha\in[0,1]$ is a credibility weight that weighs these two components. The credibility weight takes the form
$\alpha=\frac{v}{v+\kappa},$ (1.2)
where $v>0$ is a case weight indicating the amount of information contained in the individual observation $Y$ and $\kappa>0$ is the credibility coefficient (hyper-parameter) that balances between individual and collective information. This credibility coefficient $\kappa$ describes the trade-off between the different uncertainties involved at the individual and the collective risk level; see Bühlmann and Gisler (2005) for an extended discussion.
@@ 1.1.2 Transformer architectures for tabular data
The adaptation of the Transformer architecture of Vaswani et al. (2017) to tabular input data presents some challenges compared to their applications in large language models. Unlike sequential or grid-structured data, tabular features lack an inherent time-causal (positional) relationship for a canonical tensor encoding. This requires novel approaches to tokenization and positional encoding. The TabTransformer of Huang et al. (2020) pioneered the application of Transformers to tabular data by applying self-attention exclusively to categorical embeddings while processing continuous features through a separate pathway. This design choice reflects the observation that categorical variables often exhibit complex interactions that benefit from attention mechanisms, while continuous features may adequately be handled by simpler feed-forward neural network (FNN) layers.
The feature tokenizer (FT) Transformer of Gorishniy et al. (2021) extends this approach by tokenizing all features, both categorical and continuous ones, and processing them through a unified Transformer architecture. Their continuous feature tokenization uses a simple FNN embedding with an embedding space dimension being identical to the one chosen for the categorical entity embeddings. This then naturally allows one to write the input feature in tensor form, which is the suitable structure for Transformer architectures. In an actuarial context, Brauer (2024) has empirically verified the excellent performance of these FT Transformer architectures on tabular input data. Recent work on numerical feature encoding has explored various alternative approaches to handle continuous variables in Transformers. Noteworthy, piecewise linear encoding (PLE) of Gorishniy et al. (2022) provides a differentiable binning of continuous features that adapts to the data distribution.
@@ 1.1.3 In-context learning theory
ICL emerged as a surprising capability of large language models, where models can adapt to new tasks through demonstration examples provided in the input, without any parameter updates and model retraining; see Brown et al. (2020). This phenomenon has sparked intensive research into understanding the underlying mechanisms and theoretical foundations of ICL. Akyürek et al. (2023) demonstrated that linear self-attention layers can implement gradient descent on a MSE objective function when provided with appropriate in-context examples. Their analysis shows that Transformers can internally implement optimization algorithms, with attention weights encoding gradient information and value projections performing parameter updates. This finding suggests that ICL implements a form of implicit meta-learning within the forward pass. From a Bayesian perspective, Xie et al. (2022) and Müller et al. (2024) interpret ICL as implicit Bayesian inference, where the model maintains a posterior distribution over possible tasks and updates this posterior based on the context examples. Under this framework, the pre-trained distribution defines the prior, and attention mechanisms implement approximate posterior inference, providing a probabilistic foundation for understanding ICL behavior.
Recent work has identified important factors that influence ICL performance. Min et al. (2022) demonstrate that the format and distribution of demonstration examples matter more than their correctness, suggesting that ICL relies heavily on pattern matching and format recognition, while Wei et al. (2023) show that ICL capabilities emerge at scale, with larger models exhibiting qualitatively different learning behaviors than smaller ones.
Within the tabular learning context, where models are usually trained until performance on a validation loss is optimal, it would be surprising to find that ICL could lead to additional significant improvements in performance and – if this were to be the case – a new explanation for this improvement would be needed. Whereas Müller et al. (2024) have indeed shown that their TabICL model outperforms many other models, including models that only use supervised learning, as well as those using ICL, nonetheless, this could be explained by their extensive foundation model pre-training scheme using real and synthetic data. Classical supervised learning models are typically trained on a finite sample using early stopping of the gradient descent algorithm to prevent over-fitting. This early stopping may also prevent learning a more detailed, fine-grained structure, and foundation models being trained on a large body of (similar) data may be aware of such a more detailed structure.
@@ 1.1.4 Connections between credibility and meta-learning
The connection between actuarial credibility theory and modern meta-learning approaches has received limited attention despite their conceptual similarities. Both frameworks address the fundamental challenge of making predictions with limited task-specific data by leveraging broader experience. In credibility theory, this manifests as combining individual risk experience with portfolio-level information. In meta-learning, it involves leveraging experience across multiple tasks to enable rapid adaptation to new tasks. The mathematical structures underlying these approaches exhibit close parallels. The linear credibility formula (1.1) can be seen as a special case of prototype-based meta-learning, where the collective mean $\mu_{0}$ serves as a prototype and the credibility factor $\alpha$ determines the adaptation strength. Similarly, the attention mechanisms in Transformers implement a form of non-parametric regression that generalizes attention weighting to high-dimensional feature spaces. Recent work in neural processes (Garnelo et al., 2018) provides a framework that unifies these perspectives. Neural processes implement a form of meta-learning that closely resembles Bayesian inference, with the context set defining a posterior distribution over functions. This framework naturally accommodates varying amounts of context information, analogous to how credibility factors adjust (based on available experience) through the case weights $v>0$, see (1.2).
@@ 1.2 In-context learning enhanced Credibility Transformer
Our main contribution is the introduction of the In-Context Learning enhanced Credibility Transformer (ICL-Credibility Transformer), which combines the methods discussed in the previous subsections. First, the Credibility Transformer of Richman et al. (2025a) leverages the FT Transformer with a credibility mechanism that significantly enhances model training. A crucial ingredient of the Credibility Transformer is the inclusion of the CLS token of Devlin et al. (2019), that is used to learn an encoding of the input information using the classical attention layer mechanism of Vaswani et al. (2017). Instead of simply encoding and extracting this CLS token, it is credibilitized with a linear credibility mechanism that combines the feature individual CLS token encoding with a collective information token; see Richman et al. (2025a). This linear credibility combination has a very positive effect on gradient descent training, resulting in excellent predictive models. In fact, during gradient descent training, a randomized version of the credibility mechanism is applied that either selects the instance-specific CLS token or the collective information token. This randomized mechanism during model training can be interpreted as a drop-out mechanism, similar to Srivastava et al. (2014) and Wager et al. (2013), but instead of just dropping out by masking the information, the Credibility Transformer rather learns a global representation (similar to the collective mean $\mu_{0}$, see (1.1)) on the dropped out instances. This construction results in the credibilitized CLS token which is a robustified representation of the input data. The Credibility Transformer then processes these tokens through a decoder forming the predictions.
Inspired by the ICL architectures of Hollmann et al. (2023) and Müller et al. (2024) on tabular data, we modify the Credibility Transformer architecture to facilitate ICL. This raises the following architectural questions:
@@
Our primary contributions are the answers to the above questions. First, we provide an architectural innovation that enables ICL within the Credibility Transformer architecture. Second, we verify that this ICL improves the performance of the Credibility Transformer, and we give explanations for these improvements. Third, we show how the ICL-Credibility Transformer enables us to predict on new data with new categorical levels that have not been available during model training. The following items are going to be presented:
@@
    We use the same projection layer as in the base Credibility Transformer to decode the outcome decorated CLS tokens. This is justified by the fact mentioned in item (b) saying that the ICL takes place in the space of the CLS tokens, we come back to this in the training methodology below
@@
We establish formal connections between ICL mechanisms and classical credibility theory, proving that our ICL represents a credibility mechanism that regulates the impact of the context on the target instances to be predicted. We demonstrate that the attention weights serve as adaptive, data-driven credibility factors that depend on feature similarity, generalizing the fixed credibility weights of classical approaches.
@@
We deploy a two-phase training strategy that first establishes strong feature representations through a standard supervised learning method of the base Credibility Transformer. It has been verified in Richman et al. (2025a) that this learned base Credibility Transformer already presents a strong predictive model. The second training phase fine-tunes the model using ICL. During this phase, we freeze the decoder layers (i.e., the output decoder) and train the ICL mechanism, as well as the base Credibility Transformer encoder. This constrained training approach enhances the robustness of the learning process, while allowing the weights of the base Credibility Transformer encoder (which were initially optimized for supervised learning) to be modified. Keeping the output decoder frozen acts as an information bottleneck ensuring that the ICL enhanced CLS tokens remain in the same encoding space guaranteeing robustness in this second fitting step. Optionally, a third training step can be applied by jointly fine-tuning all network components, i.e., including the decoder layer; however, this requires careful regularization and a small learning rate to prevent performance degradation.
@@
We verify on the popular French Motor Third Party Liability (MTPL) dataset of Dutang et al. (2024) the outstanding performance of the ICL-Credibility Transformer. Our study quantifies the contribution of each architectural component, while explainability analyses reveal that the learned attention patterns align with actuarial intuition about risk similarity and credibility weighting. An interesting property of ICL is that we can explicitly illustrate and try to understand how individual instances in the context batch enhance the predictions; we provide empirical results explaining this case by case. A second interesting case study shows how the ICL-Credibility Transformer can be used to predict on data with new covariate levels, e.g., a new region that has not been available during model training. This generalization property (similar to transfer learning) especially draws on the context, to find the right level of integration into the existing model.